# Stopping rules for Monte Carlo methods of martingale difference type

### v1

@@ Abstract
We establish a practical and easy-to-implement sequential stopping rule for the martingale central limit theorem, focusing on Monte Carlo methods for estimating the mean of a non-iid sequence of martingale difference type. Starting with an impractical scheme based on the standard martingale central limit theorem, we progressively address its limitations from implementation perspectives in the non-asymptotic regime. Along the way, we compare the proposed schemes with their counterparts in the asymptotic regime. The developed framework has potential applications in various domains, including stochastic gradient descent methods. Numerical results are provided to demonstrate the effectiveness of the developed stopping rules in terms of reliability and complexity.
Keywords: Monte Carlo methods; martingale central limit theorem; martingale difference sequences; sequential stopping rules; stochastic gradient descent; early stopping.
@@ 1 Introduction
In Monte Carlo methods in general terms, it is a common practice to fix the total number of samples before running the simulation, while, without sufficient knowledge of the underlying distribution, it could be risky to terminate the procedure at such a priori fixed timing outright for the reason that the estimator that one is dealing with may converge extremely slowly. Moreover, the length of the simulation needs to be decided in reference to various budgetary constraints, as computing resources available are finite in practice without an exception. Those issues can be addressed by equipping the simulation procedure with a sequential (or, equivalently, adaptive) stopping rule, so that the total number of samples can be decided along the way depending on the previous computation, rather than stopping the simulation too soon or continuing the simulation endlessly.
@@
In the framework of the standard iid Monte Carlo simulation based on the central limit theorem, there has been a great deal of interest in such stopping criteria in the literature for a very long time. In the asymptotic regime, sequential stopping rules were first developed based on the first and second moments and investigated in terms of consistency and convergence rates (for instance, [3, 12, 26, 29], to name just a few). Nevertheless, in the non-asymptotic regime where the cost of computing needs to be taken into account, such asymptotics-based stopping rules may not function as theoretically expected, without extra information and restrictions, for instance, higher moments and boundedness of the underlying distribution (see, for example, [1, 9, 17] and references therein), due to an inaccurate estimate of the error probability at an early stage of the simulation procedure.
@@
Departing from the standard iid Monte Carlo simulation and its underlying standard central limit theorem, non-iid sequences have also attracted attention in the context of sequential stopping rules (for instance, [9, 10, 23]) and, more recently, machine learning. Among a variety of such non-iid sequences, the focus of the present work is set on a sequence of martingale difference type, which comes with the martingale central limit theorem. To be exact, we aim to establish a rigorous yet easy-to-implement sequential stopping rule for the martingale central limit theorem with a view towards Monte Carlo methods in estimating the mean of a sequence of random variables (just like iid Monte Carlo methods), whose law, however, keeps changing in reference to the past information within a martingale framework (unlike iid Monte Carlo methods).
@@
The relevant situation arises, for instance, from adaptive variance reduction techniques employed concurrently with the computation of the empirical mean so that the underlying randomness can be used in both procedures, such as importance sampling [21, 22], control variates and stratified sampling [32], to mention just a few. Due to their remarkable effectiveness in reducing the estimator variance without requiring much additional computing effort for running the parallel procedure, those adaptive variance reduction methods have found various fields of application, such as structural engineering [30, 31] and financial risk management [11]. The proposed framework on stopping rules for martingale difference sequences also has significant potential in machine learning optimization, particularly stochastic gradient descent methods where the stochastic gradient noises naturally form a martingale difference sequence. While conventional stopping methods often rely on validation set performance [28], our approach provides a principled alternative based on the statistical properties of the optimization trajectory itself. This is particularly of relevance given the recent variance reduction techniques for stochastic gradient descent methods [18], complementing the existing developments on validation-free stopping criteria [5].
@@
In constructing valid stopping rules of this martingale type, the primary challenge lies in the impracticality of applying the martingale central limit theorem in its standard form. Unlike the standard central limit theorem for iid Monte Carlo methods, where the limiting variance is merely unknown, the theoretical variance underpinning the martingale central limit theorem (such as in [25]) is not only unknown but also cumulative averaging. In addition, the theoretical variance here is unconditional, that is, poorly aligned with the conditional expectations inherent to martingales. Moreover, the martingale central limit theorem generally hinges on intricate and often unverifiable technical conditions, which further complicates its practical use. Given those difficulties, we explore an implementable stopping rule by progressively replacing the theoretical variance with more implementable alternatives, while carefully imposing technical conditions along the way for the corresponding stopping rules to remain valid. In particular, in its final form, the developed stopping rule operates upon a fully implementable empirical variance.
@@ 2 Preliminaries
Clearly, the standing assumption (2.1) is much weaker than the usual iid assumption and not to indicate that the sequence $\{X_{k}\}_{k\in\mathbb{N}}$ is a martingale (for which it would have to satisfy $\mathbb{E}_{k-1}[X_{k}]=X_{k-1}$ for all $k\in\mathbb{N}$ with suitable $X_{0}\in\mathcal{F}_{0}$). Instead, the assumption (2.1) can be interpreted to form a wider class than martingale difference sequences since the sequence $\{X_{k}\}_{k\in\mathbb{N}}$ is a martingale difference sequence if, moreover, $\mu=0$. It is, however, inappropriate to assume $\mu=0$ upfront or consider $Y_{k}:=X_{k}-\mu$ instead, since the constant $\mu$ is deemed unknown and is exactly what one looks to estimate in the present context. We also note that a similar assumption has been made in [9, Assumption 1.1] in the context of relative-precision stopping rules, but along with a boundedness condition on the underlying randomness (that is, $X_{n}\in[0,1]$, $a.s.$).
@@
(a) If $\{X_{k}\}_{k\in\mathbb{N}}$ is a sequence of iid random variables with finite mean $\mu$ and if $(\mathcal{F}_{k})_{k\in\mathbb{N}_{0}}$ is its natural filtration, then it satisfies Assumption 2.1. In this sense, our developments in what follows can be thought of as generalizations of the existing stopping rules for iid Monte Carlo methods. At this stage, however, we wish to emphasize that such sequences of iid random variables are outside the scope of the present study for the reason that such a way too simple situation can be treated based on much lighter conditions (for instance, [1, 3, 10, 17, 26, 29]) than we impose in what follows for the general problem setting (2.1).
@@
(c) A variety of adaptive variance reduction methods (such as [21, 22]) are also within our scope, as described in the introductory section. Those frameworks can be formulated in common as follows. Let $\{U_{k}\}_{k\in\mathbb{N}}$ be a sequence of iid uniform random vectors on $(0,1)^{d}$ and let $(\mathcal{F}_{k})_{k\in\mathbb{N}_{0}}$ be its natural filtration. For all $k\in\mathbb{N}$, we set $X_{k}=R(U_{k};{\bm{\theta}}_{k-1})$, where $R(\cdot;{\bm{\theta}})$ is an unbiased estimator for $\mu$, in the sense of $\mathbb{E}_{0}[R(U;{\bm{\theta}})]=\mu$ for $U\sim U(0,1)^{d}$, irrespective of the parameter $\bm{\theta}$}, corresponding to importance sampling and/or control variates. If the sequence $\{{\bm{\theta}}_{k}\}_{k\in\mathbb{N}_{0}}$ of the parameters is constructed in an adaptive manner to the filtration $(\mathcal{F}_{k})_{k\in\mathbb{N}}$, then it holds true that $\mathbb{E}_{k-1}[X_{k}]=\mu$, due to ${\bm{\theta}}_{k-1}\in\mathcal{F}_{k-1}$ and $U_{k}\perp\mathcal{F}_{k-1}$ for all $k\in\mathbb{N}$. The stopping rules that we develop are examined on adaptive control variates in Section 5.2. ∎
@@ 3 Problem formulation
On top of the non-assessability of the error probability, it is also a fatal drawback that the unbiasedness (Theorem 2.4 (i)) is generally no longer valid for a general $(\mathcal{F}_{m(t)})_{t\in\mathbb{N}_{0}}$-stopping time $\tau$, if the empirical batch mean is given as an output at the stopping batch (that is, $\mathbb{E}_{0}[\mu(\tau)]\neq\mu$). As an unrealistic yet illustrative example, suppose the support of the random variable $X_{k}$ under $\mathbb{P}_{0}$ is $\mathbb{R}$ for every $k\in\mathbb{N}$ and the stopping batch $\tau$ is defined as the first batch $t$ in which all elements $\{X_{k}\}_{k\in M(t)}$ are strictly positive. Then, for every $k\in M(\tau)$ in the stopping batch, the support of $X_{k}$ is $(0,+\infty)$. If $\mu=0$ (despite it is unknown in reality), then the original empirical batch mean at the stopping batch, which is now strictly positive by definition, is evidently biased for the constant $\mu$ of interest, that is, $\mathbb{E}_{0}[\mu(\tau)]>0=\mu$, $\mathbb{P}_{0}$-$a.s.$
@@
with which we compute and output the empirical batch mean $\mu_{\star}(\tau)$ of the resampled batch $(X^{\star}_{m(\tau-1)+1},\cdots,X^{\star}_{m(\tau)})$, defined by
@@
Prior to further analysis, we provide Figure 1 to illustrate the procedure for constructing the resampled empirical batch mean (3.1) along with examples below in Example 3.1.
[…]
Figure 1: Typical arrangement of batches ($m(t)$ and $M(t)$), the batched samples $\{X_{k}\}_{k\in M(t)}$, and the resampled batch mean $\mu_{\star}(\tau)$. Recall that $m(0)=0$, m(t)>m(t1)m(t)>m(t-1), $M(t)=\{m(t-1)+1,\cdots,m(t)\}$ for all $t\in\mathbb{N}$, and $|M(t)|\to+\infty$ (Assumption 2.3). Now, suppose it turns out $\tau=5$ based upon the past and latest information $\mathcal{F}_{m(5)}$. The output of our stopping rules is not the empirical batch mean $\mu(5)(=|M(5)|^{-1}\sum_{k\in M(5)}X_{k})$ along the top line, but the resampled empirical batch mean $\mu_{\star}(5)$ over the $5$th batch along the second line carrying on the past information $\mathcal{F}_{m(4)}$ (indicted by the thick arrow). The resampled sequence $\{X_{k}^{\star}\}_{k\in M(\tau)}$ is conditionally independent of the original counterpart $\{X_{k}\}_{k\in M(\tau)}$ on the stopping-time $\sigma$-field $\mathcal{F}_{m(\tau)}$.
@@
Here, we demonstrate the procedure for constructing the resampled empirical batch mean (3.1) along the concrete problem settings presented in Example 2.2. As illustrated in Figure 1, suppose it turns out $\tau=5$ along the sequence $\{X_{k}\}_{k\in\mathbb{N}}$.
(a) If $\{X_{k}\}_{k\in\mathbb{N}}$ is a sequence of iid random variables, then generate the sequence of iid random variables $\{X^{\star}_{k}\}_{k\in M(\tau)}$, each under the same law as $\mathcal{L}(X_{1})$ and independently of the original sequence $\{X_{k}\}_{k\in\mathbb{N}}$. This is exactly what is done in the existing stopping rules for the standard iid Monte Carlo simulation (for instance, [1]).
@@
(c) In the context of adaptive control variates and importance sampling, generate the sequence $\{U^{\star}_{k}\}_{k\in M(\tau)}$ of iid uniform random variables on $(0,1)$, independent of the original sequence $\{U_{k}\}_{k\in M(\tau)}$. Accordingly, in a similar manner to the construction of the augmented $\sigma$-field $\mathcal{F}^{\star}_{k}$, branch the natural filtration $(\mathcal{F}_{k})_{k\in\{0,1,\cdots,m(\tau-1)\}}$ on the basis of the resampled sequence $\{U^{\star}_{k}\}_{k\in M(\tau)}$, as well as construct $\{{\bm{\theta}}^{\star}_{k}\}_{k\in M(\tau)}$ in an adaptive manner to this branched filtration. Then, construct
@@
The motivation for introducing the resampled empirical batch mean $\mu_{\star}(\tau)$ is nothing but its unbiasedness while the one $\mu(\tau)$ based on the original sequence $\{X_{k}\}_{k\in\mathbb{N}}$ can be biased. To be exact, although the stopping batch $\tau$ is determined based on the original sequence $\{X_{k}\}_{k\in\mathbb{N}}$ (the top line of Figure 1), the output needs to be the resampled empirical batch mean $\mu_{\star}(\tau)$ at the stopping batch $\tau$ to ensure unbiasedness, as formalized below.
Theorem 3.2.

### v2

@@ Abstract
We establish a practical and easy-to-implement sequential stopping rule for the martingale central limit theorem, focusing on Monte Carlo methods for estimating the mean of a non-iid sequence of martingale difference type. Starting with an impractical scheme based on the standard martingale central limit theorem, we progressively address its limitations from implementation perspectives in the non-asymptotic regime. Along the way, we compare the proposed schemes with their counterparts in the asymptotic regime. The developed framework has potential applications in various domains. Numerical results are provided to demonstrate the effectiveness of the developed stopping rules in terms of reliability and complexity.
Keywords: Monte Carlo methods; martingale central limit theorem; martingale difference sequences; sequential stopping rules; adaptive variance reduction; early stopping.
@@ 1 Introduction
In Monte Carlo methods in general terms, it is a common practice to fix the total number of samples before running the simulation, while, without sufficient knowledge of the underlying distribution, it could be risky to terminate the procedure at such a priori fixed timing outright for the reason that the estimator that one is dealing with may converge extremely slowly. Moreover, the length of the simulation needs to be decided in reference to various budgetary constraints, as computing resources available are finite in practice without exception. Those issues can be addressed by equipping the simulation procedure with a sequential (or, equivalently, adaptive) stopping rule, so that the total number of samples can be decided along the way depending on the previous computation, rather than stopping the simulation too soon or continuing the simulation endlessly.
@@
In the framework of the standard iid Monte Carlo simulation based on the central limit theorem, there has been a great deal of interest in such stopping criteria in the literature for a very long time. In the asymptotic regime, sequential stopping rules were first developed based on the first and second moments and investigated in terms of consistency and convergence rates (for instance, [YH1965, PW1992, NS1996, RD1976], to name just a few). Nevertheless, in the non-asymptotic regime where the cost of computing needs to be taken into account, such asymptotics-based stopping rules may not function as theoretically expected, without extra information and restrictions, for instance, higher moments and boundedness of the underlying distribution (see, for example, [BHST, LWP2013, FLYA2012] and references therein), due to an inaccurate estimate of the error probability at an early stage of the simulation procedure.
@@
Departing from the standard iid Monte Carlo simulation and its underlying standard central limit theorem, non-iid sequences have also attracted attention in the context of sequential stopping rules (for instance, [LWP2013, M1968, doi:10.1287/mnsc.28.5.550, stoppingrules]) and, more recently, machine learning. Among a variety of such non-iid sequences, the focus of the present work is set on a sequence of martingale difference type, which comes with the martingale central limit theorem. To be exact, we aim to establish a rigorous yet easy-to-implement sequential stopping rule for the martingale central limit theorem with a view towards Monte Carlo methods in estimating the mean of a sequence of random variables (just like iid Monte Carlo methods), whose law, however, keeps changing in reference to the past information within a martingale framework (unlike iid Monte Carlo methods).
@@
The relevant situation arises, for instance, from adaptive variance reduction techniques employed concurrently with the computation of the empirical mean so that the underlying randomness can be used in both procedures, such as importance sampling [RK_1, RK18], control variates and stratified sampling [Kawai:2010:AOA:1734222.1734225, ISCVSS], to mention just a few. Due to their remarkable effectiveness in reducing the estimator variance without requiring much additional computing effort for running the parallel procedure, those adaptive variance reduction methods have found various fields of application, such as structural engineering [ASDIS, ReliabilitySS] and financial risk management [Glasserman2004]. The proposed framework on stopping rules for martingale difference sequences also has significant potential in machine learning optimization, particularly stochastic gradient descent methods where the stochastic gradient noises naturally form a martingale difference sequence. While conventional stopping methods often rely on validation set performance [prechelt2012], our approach provides a principled alternative based on the statistical properties of the optimization trajectory itself. This is particularly of relevance given the recent variance reduction techniques for stochastic gradient descent methods [johnson2013accelerating]. We also highlight non-asymptotic bounds for the martingale central limit theorem established for averaged stochastic gradient descent [anastasiou2019, shao2022], offering a theoretical foundation complementary to the practical stopping criteria discussed in the present study.
@@
In constructing valid stopping rules of this martingale type, the primary challenge lies in the impracticality of applying the martingale central limit theorem in its standard form. Unlike the standard central limit theorem for iid Monte Carlo methods, where the limiting variance is merely unknown, the theoretical variance underpinning the martingale central limit theorem (such as in [martingalebook2022]) is not only unknown but also cumulative averaging. In addition, the theoretical variance here is unconditional, that is, poorly aligned with the conditional expectations inherent to martingales. Moreover, the martingale central limit theorem generally hinges on intricate and often unverifiable technical conditions, which further complicates its practical use. Given those difficulties, we explore an implementable stopping rule by progressively replacing the theoretical variance with more implementable alternatives, while carefully imposing technical conditions along the way for the corresponding stopping rules to remain valid. In particular, in its final form, the developed stopping rule operates upon a fully implementable empirical variance.
@@ 2 Preliminaries
Clearly, the standing assumption (2.1) is much weaker than the usual iid assumption and not to indicate that the sequence $\{X_{k}\}_{k\in\mathbb{N}}$ is a martingale (for which it would have to satisfy $\mathbb{E}_{k-1}[X_{k}]=X_{k-1}$ for all $k\in\mathbb{N}$ with suitable $X_{0}\in\mathcal{F}_{0}$). Instead, the assumption (2.1) can be interpreted to form a wider class than martingale difference sequences since the sequence $\{X_{k}\}_{k\in\mathbb{N}}$ is a martingale difference sequence if, moreover, $\mu=0$. It is, however, inappropriate to assume $\mu=0$ upfront or consider $Y_{k}:=X_{k}-\mu$ instead, since the constant $\mu$ is deemed unknown and is exactly what one looks to estimate in the present context. We also note that a similar assumption has been made in [LWP2013, Assumption 1.1] in the context of relative-precision stopping rules, but along with a boundedness condition on the underlying randomness (that is, $X_{n}\in[0,1]$, $a.s.$).
@@
(a) If $\{X_{k}\}_{k\in\mathbb{N}}$ is a sequence of iid random variables with finite mean $\mu$ and if $(\mathcal{F}_{k})_{k\in\mathbb{N}_{0}}$ is its natural filtration, then it satisfies Assumption 2.1. In this sense, our developments in what follows can be thought of as generalizations of the existing stopping rules for iid Monte Carlo methods. At this stage, however, we wish to emphasize that such sequences of iid random variables are outside the scope of the present study for the reason that such a way too simple situation can be treated based on much lighter conditions (for instance, [BHST, YH1965, M1968, FLYA2012, NS1996, RD1976]) than we impose in what follows for the general problem setting (2.1).
@@
(c) A variety of adaptive variance reduction methods (such as [RK_1, RK18]) are also within our scope, as described in the introductory section. Those frameworks can be formulated in common as follows. Let $\{U_{k}\}_{k\in\mathbb{N}}$ be a sequence of iid uniform random vectors on $(0,1)^{d}$ and let $(\mathcal{F}_{k})_{k\in\mathbb{N}_{0}}$ be its natural filtration. For all $k\in\mathbb{N}$, we set $X_{k}=R(U_{k};{\bm{\theta}}_{k-1})$, where $R(\cdot;{\bm{\theta}})$ is an unbiased estimator for $\mu$, in the sense of $\mathbb{E}_{0}[R(U;{\bm{\theta}})]=\mu$ for $U\sim U(0,1)^{d}$, irrespective of the parameter $\bm{\theta}$}, corresponding to importance sampling and/or control variates. If the sequence $\{{\bm{\theta}}_{k}\}_{k\in\mathbb{N}_{0}}$ of the parameters is constructed in an adaptive manner to the filtration $(\mathcal{F}_{k})_{k\in\mathbb{N}}$, then it holds true that $\mathbb{E}_{k-1}[X_{k}]=\mu$, due to ${\bm{\theta}}_{k-1}\in\mathcal{F}_{k-1}$ and $U_{k}\perp\mathcal{F}_{k-1}$ for all $k\in\mathbb{N}$. The stopping rules that we develop are examined on adaptive control variates in Section 5.2. ∎
@@ 3 Problem formulation
On top of the non-assessability of the error probability, it is also a crucial drawback that the unbiasedness (Theorem 2.4 (i)) is generally no longer valid for a general $(\mathcal{F}_{m(t)})_{t\in\mathbb{N}_{0}}$-stopping time $\tau$, if the empirical batch mean is given as an output at the stopping batch (that is, $\mathbb{E}_{0}[\mu(\tau)]\neq\mu$). As an unrealistic yet illustrative example, suppose the support of the random variable $X_{k}$ under $\mathbb{P}_{0}$ is $\mathbb{R}$ for every $k\in\mathbb{N}$ and the stopping batch $\tau$ is defined as the first batch $t$ in which all elements $\{X_{k}\}_{k\in M(t)}$ are strictly positive. Then, for every $k\in M(\tau)$ in the stopping batch, the support of $X_{k}$ is $(0,+\infty)$. If $\mu=0$ (despite it is unknown in reality), then the original empirical batch mean at the stopping batch, which is now strictly positive by definition, is evidently biased for the constant $\mu$ of interest, that is, $\mathbb{E}_{0}[\mu(\tau)]>0=\mu$, $\mathbb{P}_{0}$-$a.s.$
@@
with which we compute and output the empirical batch mean $\mu_{\star}(\tau)$ of the regenerated batch $(X^{\star}_{m(\tau-1)+1},\cdots,X^{\star}_{m(\tau)})$, defined by
@@
For illustrative purposes, we provide Figure 1 to illustrate the procedure for constructing the regenerated empirical batch mean (3.1).
[…]
Figure 1: Typical arrangement of batches ($m(t)$ and $M(t)$), the batched samples $\{X_{k}\}_{k\in M(t)}$, and the regenerated batch mean $\mu_{\star}(\tau)$. Recall that $m(0)=0$, m(t)>m(t1)m(t)>m(t-1), $M(t)=\{m(t-1)+1,\cdots,m(t)\}$ for all $t\in\mathbb{N}$, and $|M(t)|\to+\infty$ (Assumption 2.3). Now, suppose it turns out $\tau=5$ based upon the past and latest information $\mathcal{F}_{m(5)}$. The output of our stopping rules is not the empirical batch mean $\mu(5)(=|M(5)|^{-1}\sum_{k\in M(5)}X_{k})$ along the top line, but the regenerated empirical batch mean $\mu_{\star}(5)$ over the $5$th batch along the second line carrying on the past information $\mathcal{F}_{m(4)}$ (indicated by the thick arrow). The regenerated sequence $\{X_{k}^{\star}\}_{k\in M(\tau)}$ is conditionally independent of the original counterpart $\{X_{k}\}_{k\in M(\tau)}$ on the stopping-time $\sigma$-field $\mathcal{F}_{m(\tau)}$.
The motivation for introducing the regenerated empirical batch mean $\mu_{\star}(\tau)$ is nothing but its unbiasedness while the one $\mu(\tau)$ based on the original sequence $\{X_{k}\}_{k\in\mathbb{N}}$ can be biased. To be exact, although the stopping batch $\tau$ is determined based on the original sequence $\{X_{k}\}_{k\in\mathbb{N}}$ (the top line of Figure 1), the output needs to be the regenerated empirical batch mean $\mu_{\star}(\tau)$ at the stopping batch $\tau$ to ensure unbiasedness, as formalized below.
Theorem 3.1.
@@
Before closing this section, we provide a more detailed illustration of how to construct the regenerated empirical batch mean (3.1), as depicted in Figure 1, based upon the three problem settings already outlined in Example 2.2.
@@
Here, we demonstrate the procedure for constructing the regenerated empirical batch mean (3.1) along the concrete problem settings presented in Example 2.2. As illustrated in Figure 1, suppose it turns out $\tau=5$ along the sequence $\{X_{k}\}_{k\in\mathbb{N}}$.
(a) If $\{X_{k}\}_{k\in\mathbb{N}}$ is a sequence of iid random variables, then generate the sequence of iid random variables $\{X^{\star}_{k}\}_{k\in M(\tau)}$, each under the same law as $\mathcal{L}(X_{1})$ and independently of the original sequence $\{X_{k}\}_{k\in\mathbb{N}}$. This is exactly what is done in the existing stopping rules for the standard iid Monte Carlo simulation (for instance, [BHST]).
@@
(c) In the context of adaptive control variates and importance sampling, generate the sequence $\{U^{\star}_{k}\}_{k\in M(\tau)}$ of iid uniform random variables on $(0,1)$, independent of the original sequence $\{U_{k}\}_{k\in M(\tau)}$. Accordingly, in a similar manner to the construction of the augmented $\sigma$-field $\mathcal{F}^{\star}_{k}$, branch the natural filtration $(\mathcal{F}_{k})_{k\in\{0,1,\cdots,m(\tau-1)\}}$ on the basis of the regenerated sequence $\{U^{\star}_{k}\}_{k\in M(\tau)}$, as well as construct $\{{\bm{\theta}}^{\star}_{k}\}_{k\in M(\tau)}$ in an adaptive manner to this branched filtration. Then, construct